# Retroreflective surface optimisation for optical cavities with custom mirror profiles

### v1

@@ Appendix D Derivation of critical diameter, separation, and best geometry for coupling to two emitters
Starting with the derivation of $D_{\mathrm{crit}}$, we are concerned with the minimum possible clipping loss for any Gaussian mode, given the constraints of $L$ and $D$, which is obtained when $w_{m}$ is minimised. We aim to find the critical diameter $D=D_{\mathrm{crit}}$ such that the clipping loss $\mathcal{L}_{\mathrm{clip}}$ and non-clipping loss $\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0}$ balance. This is because, if $\mathcal{L}_{\mathrm{clip}}$ dominates, the total cavity losses are high, and the performance low for any spherical mirror, whereas if $\mathcal{L}_{\mathrm{clip}}$ is negligible, it is possible to achieve loss at the minimum level of $\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0}$, and thus achieve high cavity performance. The critical diameter achieving this balance is therefore
$D_{\mathrm{crit}}=\sqrt{\frac{\lambda L}{\pi}}\sqrt{2\left(\log(2)-\log\left(\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0}\right)\right)},$ (32)
obtained through rearrangement of Eq. 30 with $w_{m}=\sqrt{\lambda L/\pi}$ as the minimum waist at mirror from Eq. 31.
We now derive the critical separation $s_{\mathrm{crit}}$, defined such that the mode that has the minimum waist size at both emitters separated by the critical separation $s_{\mathrm{crit}}$, must have clipping loss on the mirror $\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0}$. We know (Eq. 31) that the smallest beam waists in the emitter planes (separated by $s$), are found with central waist $w_{0}=\sqrt{\lambda s/(2\pi)}$. This leads to a waist at the mirror of
@@
where the approximation that $L^{2}\gg s^{2}$ has been made in the right hand expression for algebraic convenience. This approximation is often reasonable, as the emitter separation could be far smaller than the cavity length, but not bigger. However, we would expect to see some discrepancies when $s\sim L$. Taking this assumption, substituting $w_{m}$ into Eq. 30 where the clipping loss should equal the non-clipping internal loss $(\mathcal{L}_{\mathrm{clip}}=\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0})$, and rearranging, yields
$s_{\mathrm{crit}}=\frac{\lambda L^{2}}{\pi D^{2}}\left\{\log(2)-\log\left(\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0}\right)\right\}.$ (34)
Finally, we discuss the large mirror diameter $D$ limit to $C_{\mathrm{int}}$ when coupling to two emitters separated by $s$. As mirror diameter is large, the clipping losses will be negligible, so our only concern is minimising the waist on both emitters. From Eq. 31, we know the minimum waist at the emitters is $w_{e}=\sqrt{\lambda s/\pi}$.
@@
Additionally, as the clipping losses are expected to be negligible, we expect $\mathcal{L}_{\mathrm{int}}=\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0}$. Applying Eq. 10 of the main text results in an internal cooperativity of
$C_{\mathrm{int}}=\frac{6\lambda}{\pi s\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0}}.$ (36)

### v2

@@ Appendix E Derivation of critical diameter, separation, and best geometry for coupling to two emitters
Starting with the derivation of $D_{\mathrm{crit}}$, we are concerned with the minimum possible clipping loss for any Gaussian mode, given the constraints of $L$ and $D$, which is obtained when $w_{m}$ is minimised. We aim to find the critical diameter $D=D_{\mathrm{crit}}$ such that the clipping loss $\mathcal{L}_{\mathrm{clip}}$ and non-clipping loss $\mathcal{L}_{\mathrm{abs}}$ balance. This is because, if $\mathcal{L}_{\mathrm{clip}}$ dominates, the total cavity losses are high, and the performance low for any spherical mirror, whereas if $\mathcal{L}_{\mathrm{clip}}$ is negligible, it is possible to achieve loss at the minimum level of $\mathcal{L}_{\mathrm{abs}}$, and thus achieve high cavity performance. The critical diameter achieving this balance is therefore
$D_{\mathrm{crit}}=\sqrt{\frac{\lambda L}{\pi}}\sqrt{2\left(\log(2)-\log\left(\mathcal{L}_{\mathrm{abs}}\right)\right)},$ (40)
obtained through rearrangement of Eq. (38) with $w_{m}=\sqrt{\lambda L/\pi}$ as the minimum waist at mirror from Eq. (39).
We now derive the critical separation $s_{\mathrm{crit}}$, defined such that the mode that has the minimum waist size at both emitters separated by the critical separation $s_{\mathrm{crit}}$, must have clipping loss on the mirror $\mathcal{L}_{\mathrm{abs}}$. We know (Eq. (39)) that the smallest beam waists in the emitter planes (separated by $s$), are found with central waist $w_{0}=\sqrt{\lambda s/(2\pi)}$. This leads to a waist at the mirror of
@@
where the approximation that $L^{2}\gg s^{2}$ has been made in the right hand expression for algebraic convenience. This approximation is often reasonable, as the emitter separation could be far smaller than the cavity length, but not bigger. However, we would expect to see some discrepancies when $s\sim L$. Taking this assumption, substituting $w_{m}$ into Eq. (38) where the clipping loss should equal the non-clipping internal loss $(\mathcal{L}_{\mathrm{clip}}=\mathcal{L}_{\mathrm{abs}})$, and rearranging, yields
$s_{\mathrm{crit}}=\frac{\lambda L^{2}}{\pi D^{2}}\left\{\log(2)-\log\left(\mathcal{L}_{\mathrm{abs}}\right)\right\}.$ (42)
Finally, we discuss the large mirror diameter $D$ limit to $C_{\mathrm{int}}$ when coupling to two emitters separated by $s$. As mirror diameter is large, the clipping losses will be negligible, so our only concern is minimising the waist on both emitters. From Eq. (39), we know the minimum waist at the emitters is $w_{e}=\sqrt{\lambda s/\pi}$.
@@
Additionally, as the clipping losses are expected to be negligible, we expect $\mathcal{L}_{\mathrm{int}}=\mathcal{L}_{\mathrm{abs}}$. Applying Eq. (11) of the main text results in an internal cooperativity of
$C_{\mathrm{int}}=\frac{6\lambda}{\pi s\mathcal{L}_{\mathrm{abs}}}.$ (44)
Appendix F Numerical parameters for simulations
The data presented in the main text compares retroreflective optimised cavities against the cavities with the best possible spherical/circular mirrors for the same geometric constraints. This procedure has three steps. The first is to scan the possible spherical/circular mirrors in order to find both the best performance with spherical mirrors, and the basis functions to use for retroreflective optimisation. These calculations were performed using mode mixing using $N_{\mathrm{sph.}}$ basis states and numerical integration with $s_{\mathrm{sph.}}$ samples in the Riemann sum (linearly spaced in either $r$ or $x$ coordinate). The second step is the first component of the retroreflective optimisation i.e. the determination of $\epsilon_{\mathrm{T}}$. This optimises the chosen metric on the assumption of perfect retroreflection, where the chosen metric is evaluated in a basis of size $N$, of which we choose to optimise the coefficients of the first $N_{\mathrm{opt.}}$ basis states, and leave the remainder zero. The third step is to take the optimised $\epsilon_{\mathrm{T}}$, and use the surface reconstruction method (Sec. III main text), where the surface is initially sampled (again, linearly in $r$ or $x$) using $s$ small elements, and mode mixing is performed in a basis of size $N$. A summary of the parameters used for the simulations can be found in Table 1. From this, we highlight that the optimisation for Fig. 3, where the emitter was displaced towards the edge of the cavity, required a very large number of samples for its mode mixing integrals to converge, but such sampling was not necessary for cases with the emitter closer to the cavity centre.
Example $N_{\mathrm{sph.}}$ $s_{\mathrm{sph.}}$ NN $N_{\mathrm{opt.}}$ $s$
Fig. 1 171 15000 171 136 15000
Fig. 2 171 15000 171 136 15000
Fig. 3 50 300 250 200 500000
Fig. 5 examples 171 15000 171 136 15000
Fig. 5 scan 171 5000 171 136 5000
Fig. 6 examples 171 15000 171 136 15000
Fig. 6 scan 111 5000 111 81 5000
Table 1: Parameters used for the retroreflective optimisations conducted in the main text. The columns indicate the parameter, and the rows discuss the figure, or part of the figure, where the relevant optimisation is used.
The second part of this procedure (i.e. the first part of the retroreflective optimisation method) optimises $\epsilon_{\mathrm{T}}$ by changing the coefficients of $N_{\mathrm{opt.}}$ states in a basis of size $N$. The number of coefficients ultimately corresponds to the number of degrees of freedom in the target mode, and consequently the resultant mirror surfaces. If $N_{\mathrm{opt.}}$ is too low, there will be not enough freedom to describe the optimum $\epsilon_{\mathrm{T}}$, and thus the method will not report the optimum performance. Data indicating how the retroreflective performance (i.e. the cavity performance assuming perfect retroreflection) varies with $N_{\mathrm{opt.}}$ and $N$ are shown in Fig. 10. Here we see that, for most examples, $\leq 50$ basis states and optimised coefficients are sufficient to find the optimised performance to a high degree of accuracy.
Figure 10: The dependence of the cavity performance assuming perfect retroreflection on the basis size $N$ and number of variable coefficients $N_{\mathrm{opt.}}$, where $N_{\mathrm{opt.}}\leq N$. Note that, because this studies the convergence of the first part of retroreflective optimisation only, no mode mixing calculation is conducted. Each panel shows the convergence data for the titled figure. The black band on the colour bar indicates the retroreflected expected performance for the best spherical mirror. The cyan band in the colour bar and as a contour on the performance map indicates the boundary above which the reported performance has increased by 90% of the total increase from the spherical case to the optimised value reported for the largest basis sizes in that panel’s data. The red cross indicates the parameters chosen for the case study examples in each figure. Note that the black band in panel d) is very close to zero, because the best spherical mirror cavity produced a highly mixed mode, where the nominal fundamental mode was strongly clipped.